### v1

# Exploring the Landscape of Non-Equilibrium Memories with Neural Cellular Automata

Abstract
We investigate the landscape of many-body memories: families of local non-equilibrium dynamics that retain information about their initial conditions for thermodynamically long time scales, even in the presence of arbitrary perturbations. In two dimensions, the only well-studied memory is Toom’s rule. Using a combination of rigorous proofs and machine learning methods, we show that the landscape of 2D memories is in fact quite vast. We discover memories that correct errors in ways qualitatively distinct from Toom’s rule, have ordered phases stabilized by fluctuations, and preserve information only in the presence of noise. Taken together, our results show that physical systems can perform robust information storage in many distinct ways, and demonstrate that the physics of many-body memories is richer than previously realized. Interactive visualizations of the dynamics studied in this work are available at https://memorynca.github.io/2D.
Introduction:
The design of many-body systems that retain memory of their initial conditions in the presence of perturbations, i.e. systems which function as robust many-body memories, is a fundamental problem at the intersection of non-equilibrium statistical mechanics [1], fault-tolerant computation [2], and biological information processing [3]. Memory in locally-interacting systems must be achieved through collective, distributed error correction: errors must be autonomously detected and corrected by the noisy dynamics itself, and the general mechanisms by which this may be accomplished are only partially understood [4, 2, 5, 6, 7, 8, 9, 10, 11, 12, 13].
This work investigates non-equilibrium many-body memories in 2D spin systems. The dynamics we study are designed to remember a single bit of information, which they encode in the sign of the system’s magnetization. Concretely, we consider dynamics acting on spins $s_{\mathbf{r}}\in\{\pm 1\}$ located on the sites of an $L\times L$ lattice with periodic boundary conditions. The dynamics is defined as a probabilistic Cellular Automaton (CA), specified by a noise strength $p\in[0,1]$, a noise bias $\eta\in[-1,1]$, and a CA rule $\mathcal{A}$. Letting $\mathcal{R}_{\mathbf{r}}$ denote the $3\times 3$ block of sites centered at a site $\mathbf{r}$, the dynamics updates $s_{\mathbf{r}}$ according to a particular function $\mathcal{A}$ of the spins on $\mathcal{R}_{\mathbf{r}}$. When noise is included, the dynamics acts by sending $s_{\mathbf{r}}\mapsto\mathcal{A}(s|_{\mathcal{R}_{\mathbf{r}}})$ with probability 1p1-p, and $s_{\mathbf{r}}\mapsto\pm 1$ with probability $p(1\pm\eta)/2$.
In the math and computer science literature, CA are normally updated synchronously: they are run in discrete time, and all sites are updated at once. From the physics perspective, an asynchronous update scheme—where the sites update at times governed by independent Poisson processes—is often more natural, as it corresponds to a continuous-time Markov process. We will consider both types of updates in this work.
We say that $\mathcal{A}$ defines a robust memory if there is an open ball in the $(p,\eta)$ plane within which the dynamics preserves the sign of the magnetization $m=L^{-2}\sum_{\mathbf{r}}s_{\mathbf{r}}$ for a time that diverges as $L\rightarrow\infty$. Having a robust memory is only possible if regions of errors are corrected in a time proportional to their linear size: this is because a bias $\eta$ favoring a minority domain causes it to ballistically expand, and this expansion rapidly destroys the encoded information if the minority domain is not quickly eroded by the dynamics. In this context, the problem of constructing a robust memory is therefore equivalent to finding a CA that quickly identifies and corrects arbitrarily-shaped domains of errors  [^1] An example of dynamics which does not erase minority domains quickly enough is Glauber dynamics for the 2D Ising model at $T<T_{c}$. A domain of size $R$ is corrected only in time $\sim R^{2}$, which leads to the memory being destroyed in the presence of biased noise (or essentially equivalently, in the presence of a small magnetic field). .
[…]
Figure 1: Histories showing the correction of large minority domains of $1$ spins (black) for different memories, with the noise bias $\eta=+1$ chosen to maximally favor the error. Time runs left to right. b) A schematic phase diagram in the $(p,\eta)$ plane for typical rules found with our NCA architecture. c) Schematic phase diagram of the $\mathsf{D}$ rule, with a re-entrant noise-stabilized memory phase; the tip of the memory phase is around $(p,\eta)\approx(0.4,0.08)$. $\sf d)$} $p_{\sf log}$} against $p$ for asynchronous updates, shown for the $\mathsf{A}$ automaton at $\eta=0$. The dashed lines are fits to ${p_{\sf log}}\propto p^{\lambda L}$ with $\lambda\approx 1.3$. $\sf e)$} the same as $\sf d$}, but for the $\mathsf{C}$} automaton at $\eta=-1$; here $\lambda\approx 2.6$.
A first class of examples:
as a first exact result, we prove that the triangular geometry of Toom’s rule is in fact not special, and that most ways of doing local majority voting yield robust memories under synchronous updates:
Theorem 1 (asymmetric majority voters are robust).
Suppose $\mathcal{A}(s|_{\mathcal{R}_{\mathbf{r}}})={\sf maj}(\{s_{\mathbf{r}^{\prime}\in\mathcal{V}_{\mathbf{r}}}\})$ performs a majority vote on a subset of spins in the neighborhood $\mathcal{V}_{\mathbf{r}}\subset\mathcal{R}_{\mathbf{r}}$. Then $\mathcal{A}$ is a robust memory under synchronous updates iff $\mathcal{V}_{\mathbf{r}}$ is not invariant under a $\pi$ rotation $C_{\pi}$ about any lattice point.
The proof makes use of a theorem of Toom [6] regarding monotone CA (viz. those for which $\mathcal{A}$ defines a monotonic Boolean function  [^2] A function $f(x)\,:\,\{-1,1\}^{n}\rightarrow\{-1,1\}$ is monotonic if increasing the number of 1s in the input cannot change an output of 1 to an output of 1-1., such as a majority vote), which states that any monotone eroder is a robust memory at small enough noise strengths. We prove erosion by proving that minority domains are always eroded ballistically along certain directions if $\mathcal{V}_{\mathbf{r}}$ has no $C_{\pi}$ symmetry. Conversely, with $C_{\pi}$ symmetry, we show that there are always certain shapes of minority domains that, regardless of their size, are never corrected by $\mathcal{A}$; under biased noise, these domains grow and rapidly destroy the encoded information. The details are provided in App. III.
As a concrete example, consider the CA rule $\mathsf{A}$ defined by the voting region $\mathcal{V}_{\mathbf{r}}=\mathbf{r}+\{{\mathbf{\hat{y}}},{\mathbf{\hat{x}}}+{\mathbf{\hat{y}}},{\mathbf{\hat{x}}},-{\mathbf{\hat{y}}},-{\mathbf{\hat{x}}}\}$. As illustrated in the top row of Fig. 1 a, $\mathsf{A}$ erodes minority domains by mapping them to a series of increasingly small pengaton-shaped regions.
As stated, theorem 1 only applies to dynamics with synchronous updates. The extension to asynchronous updates is nontrivial, as there exist systems which retain memory only when updated synchronously [18]. To establish results for asynchronous dynamics, we therefore turn to numerics. We use Monte Carlo simulations to estimate the logical failure rate, defined as the probability for a magnetization flip to occur by time $O(L)$ when the system is initialized in its least stable logical state:
$p_{\sf log}=\max_{s}\mathsf{P}[{\sf maj}(\mathcal{A}^{cL}(s))\neq{\sf maj}(s)],$ (1)
where the maximum is over states $s$ with long-lived magnetizations (e.g. the states with all $1$s or all 1-1s), and $c$ is a constant, fixed at $c=5$ in our numerics. For any memory we must have $\lim_{L\rightarrow\infty}p_{\sf log}=0$, and for the CA under consideration we expect, for small enough $p$, the scaling
$p_{\sf log}\sim(p/p_{c})^{\lambda L}$ (2)
for constants $p_{c},\lambda$  [^3] This is due to the fact that for all of the memories considered in the main text, flipping $\Theta(2L)$ spins along two non-contractible loops that span each cycle of the torus is sufficient to cause a magnetization flip.. $p_{\sf log}$} for $\mathsf{A}$ dynamics with asynchronous updates and $\eta=0$ is plotted in Fig. 1 d, where we observe a threshold crossing at $p_{c}\approx 26\%$—significantly higher than the $p_{c,{\sf Toom}}\sim 14\%$ of Toom’s rule [20, 10].
Neural cellular automata:
Do there exist robust memories beyond asymmetric majority voters? Addressing this question via brute-force search is impossible: the number of CA rules taking $3\times 3$ blocks of spins as input is $2^{2^{9}}$—vastly greater than the number of atoms in the observable universe—and random rules are empirically observed to have a negligibly small chance of being robust memories. To more efficiently explore the space of possible dynamics, we use a Machine Learning (ML)-guided approach inspired by the Neural Cellular Automaton (NCA) framework [8]. In this setup, $\mathcal{A}$ is initially parametrized as a random continuous-state CA rule (operating on $\mathbb{R}$-valued spins), implemented by a small neural network. The neural network is optimized using gradient descent to minimize a loss function that is designed to measure $\mathcal{A}$’s ability to erase large domains of errors in the presence of noise. Including large error regions in the training data, instead of relying on their spontaneous occurrence from noisy dynamics, enables faster and more targeted learning. After a fixed number of optimization steps, we discretize $\mathcal{A}$ to $\mathbb{Z}_{2}$-valued spins. Further details on the NCA architecture and training are provided in App. II.
In a set of 1000 training runs with different random seeds, 37 cellular automata converged to robust memories. None of these 37 memory rules were symmetric under $s_{\mathbf{r}}\mapsto-s_{\mathbf{r}}$ spin-flip, and all but 6 were monotonic. Symmetric rules occurred only outside the memory class, where they were common: out of the 387 symmetric rules observed, all were monotonic, and none were memories. Monotonicity was similarly widespread beyond the memory class, with 994 monotonic rules in total. In most cases where the learned rule was not a memory, the dynamics was observed to perform something similar to a $C_{\pi}$-symmetric majority vote. Finally, while training was performed only with synchronous updates, all of the learned memory rules were observed to remain robust after making the updates asynchronous.
Representative examples of learned memories are shown in the last three rows of Fig. 1 a. The automaton $\mathsf{B}$ in the second row erodes domains by shrinking them into small triangle shapes, while $\mathsf{C}$ in the third row squeezes domains along an oblique angle. Both $\mathsf{B},\mathsf{C}$} are monotonic, and possess no symmetries. These rules have phase diagrams and a scaling of $p_{\sf log}$} qualitatively similar to that of $\mathsf{A}$, and their thresholds are tabulated in Tab. 1.
$p_{c}$ (%) $\sf A$} $\sf B$} $\sf C$} $\sf D$}
async. 4/26/4 4/22/3 8/26/10 noise stab.
sync. 5/32/5 6/32/5 9/34/11 noise stab.
Table 1: Approximate thresholds (in percent) of the memories studied in the main text, for both asynchronous (top) and synchronous (bottom) updates. The three values $x/y/z$ indicate the maximum noise probability thresholds at biases of $\eta=1,0,$ and 1-1, respectively. “noise-stab” indicates that the memory is present only at nonzero noise; see the main text.
Noise-stabilized order:
The automaton $\mathsf{D}$ shown in the fourth row of Fig. 1 a, which is one of the non-monotonic rules found in training, is rather unusual. At $p=0$ it has exponentially many absorbing states, created by inserting small islands of 1-1s in a background of $1$s (in the language of [9, 21], it is thus not a linear eroder). Under unbiased noise these islands percolate, resulting in a unique steady state dominated by 1-1s. With a modest amount of noise biased to favor $1$, however, the noise and CA dynamics render the islands unstable. This results in the phase diagram of Fig. 1 c, featuring a “re-entrant” memory phase in an intermediate range of $p$. The $\mathsf{D}$ automaton is thus a noise-stabilized memory, and its existence demonstrates that even non-eroding automata can define robust memories. This phenomenon can in some sense be regarded as a non-equilibrium analogue of the Pomeranchuk effect [22, 23], but arises due to a quite different microscopic mechanism. The only other example of noise-stabilized order the authors are aware of is the recent Ref. [24], which is based on a modification of Gacs’ complex 1D automaton [2].
The failure of mean-field theory:
The most natural first pass at theoretically understanding the phase diagrams observed so far is to perform a mean-field (MF) analysis. This is done by writing down a master equation for $\partial_{t}\langle s_{\mathbf{r}}\rangle$ (with the expectation value $\langle\cdot\rangle$ taken in the non-equilibrium steady state), and assuming that correlation functions factorize as $\langle s_{\mathbf{r}}s_{\mathbf{r}^{\prime}}\rangle\approx\langle s_{\mathbf{r}}\rangle\langle s_{\mathbf{r}^{\prime}}\rangle\equiv m^{2}$, where we have assumed $\langle s_{\mathbf{r}}\rangle$ is independent of $\mathbf{r}$ (as it is for all of the steady states we consider). Doing this yields (see App. IV)
$\partial_{t}m=p\eta+(1-p)\mathcal{A}_{MF}(m)-m\equiv-\partial F_{\rm eff}/\partial m$ (3)
where $\mathcal{A}_{MF}(m)$ is obtained by writing down $\mathcal{A}(s|_{\mathcal{R}_{\mathbf{r}}})$ as a polynomial in the spins on $\mathcal{R}_{\mathbf{r}}$, and substituting $s_{\mathbf{r}}\mapsto m$.
MF theory predicts an ordered phase when the effective free energy $F_{\rm eff}$ develops multiple distinct minima. For memories where $\mathcal{A}$ performs a majority vote, $F_{\rm eff}$ is a conventional-looking free energy with multiple minima at $p<p_{c,MF}$, and for the $\mathsf{A}$ automaton we obtain $p_{c,MF}=7/15>p_{c}$. Having $p_{c,MF}>p_{c}$ is of course unsurprising: since MF theory neglects fluctuations, it is expected to always overestimate the tendency towards ordering.
Interestingly, the CA rules $\mathsf{B},\mathsf{C},\mathsf{D}$} found in training are observed to be unstable in MF: for these rules, $F_{\rm eff}$ always has a unique minimum, regardless of $p$. These are the first examples known to the authors of systems where MF theory entirely misses an ordered phase, and incorrectly predicts a disordered phase for all $p$. The order in these systems is thus fluctuation-stabilized  [^4] This phenomenon might be called “order by disorder”, but is quite distinct from what is usually meant by this phrase, which conventionally refers to situations where (fine-tuned) systems with many degenerate ground states have their degeneracy lifted by fluctuations. Conventional “order by disorder” decreases the number of degenerate states, and thus in some sense actually reduces the amount of order. By contrast, the dynamics considered in this work uses fluctuations to increase the number of steady states, thereby increasing the amount of order. , and future work [18] will show how fluctuations beyond MF stabilize the ordered phase.
Synchronicity transitions: By explicitly modifying their rule tables, the learned rules can be used as jumping-off points for exploring different types of error-correcting dynamics. As one illustrative example, we discuss a modification that changes how these automata behave with respect to changes in synchronicity.
Figure 2: Synchronicity transitions in the modified CA rules $\mathsf{A}^{\prime},\mathsf{B}^{\prime},\mathsf{C}^{\prime},\mathsf{D}^{\prime}$. Dark colors are 11s, and each image shows the state of a system after 50 steps of noiseless dynamics following a quench from the all-(1)(-1)s state. The number at bottom right of each panel indicates the value of $\alpha$, with $\alpha_{c}$ the critical value of the synchronicity parameter. For rules $\mathsf{A}^{\prime},\mathsf{B}^{\prime},\mathsf{C}^{\prime}$, the all-1s state is rapidly reached when $\alpha>\alpha_{c}$. For $\mathsf{D}^{\prime}$, the patterns shown in the bottom-right panel persist for infinite times in the absence of noise.
In the absence of noise, all of the learned NCA rules have the all 1s and all 1-1s configurations as absorbing states. Consider destroying the 1-1 absorbing state by modifying the CA rule to send $\mathcal{A}(s|_{\mathcal{R}_{\mathbf{r}}})\mapsto+1$ if all spins in $\mathcal{R}_{\mathbf{r}}$ are 1-1. For the examples considered above, this change immediately destroys the memory for synchronous updates, even in the absence of noise: an initial state with mostly 1-1s is immediately mapped to a state with mostly $1$s, leading to an irrecoverable loss of the encoded information. On the other hand, when the updates are asynchronous, isolated errors introduced by the modification to the transition rule have a high probability of being corrected by updates on adjacent sites, and the memory is retained up to a non-zero (but smaller) critical noise strength. These automata thus define asynchronicity-protected memories, and stand in contrast with all other known memory CA, which become more robust when the updates are made synchronous.
To further explore this phenomenon, consider updates where a fraction $\alpha\in[0,1]$ of sites are simultaneously updated at each time step. At small enough $p$, each of the studied memories undergoes a percolation transition at an intermediate update fraction $\alpha_{c}$ between a robust memory phase (at small $\alpha$), and a trivial phase dominated by $1$s (at large $\alpha$); this phenomenon is somewhat similar to the synchronicity transitions that occur in other (non-robust) CA, such as Conway’s game of life [26, 27]. As shown in Fig. 2, the spatial spin patterns produced across this transition can be quite rich.
Figure 3: Top: Probability that a rule remains a memory $P_{\sf mem}$ as a function of the number of “mutations” made to its rule table; each data point is based on sampling 5000 randomly mutated rules. Bottom: Hamming distances between the rule tables of the learned NCA and majority-vote memories.
The rule landscape:
The modification just discussed was chosen in a quite special way, and it is natural ask what happens when random modifications are made. To study this, note that a given CA is determined by a binary string of length $2^{9}$, with each digit specifying the output of the CA rule on one of the $2^{9}$ possible inputs. We then take this binary string, which may be regarded as that CA’s “genetic code”, and subject it to “mutations” by randomly flipping bits. In the top panel of Fig. 3, we plot the probability $P_{\sf mem}$ that the studied rules remain a memory as a function of the number of inflicted mutations, where $P_{\sf mem}$ is estimated as the fraction of mutated rules which remain able to shrink both types of minority domains in the absence of noise. Rules $\mathsf{A},\mathsf{B},\mathsf{C}$} are seen to be quite robust to mutations, with $\gtrsim 50$ mutations required to bring $P_{\sf mem}$ below $0.5$.
This result shows that each learned CA rule carves out an extended region of robust memories within the space of all CA rules. How are the regions associated to different rules connected with one another? We address this by computing the Hamming distances between the rule tables of each learned NCA rule, as well as the distances between these rules and those defined by asymmetric majority votes. The results in the bottom panel of Fig. 3 show that the NCA rules lie in a bubble of Hamming radius $\sim 50$, and appear to be distinctly separated from the rules based on majority votes. Fleshing out the topology of this landscape is left to future work.
Outlook:
The phenomena discovered in this work indicate that the physics of non-equilibrium memories is quite diverse, and far from fully understood. Immediate future reserach directions include generalizing the NCA framework to $\mathbb{Z}_{N}$ and $U(1)$ memories, investigating the phase transitions that occur when the memory is lost, and theoretically characterizing the distinct universality classes of local error-correcting dynamics.
We close with some rather speculative remarks. The systems studied in this work are capable of “homeostasis”: the robust protection of (here, a single bit of) information against degredation by noise. It is natural to wonder about mechanisms by which richer behavior, like nontrivial information processing and environmental adaptation, can robustly emerge in locally-interacting systems. It is well known that simple CA rules can display remarkably life-like emergent behavior [28, 29, 8], but in none of these cases has the problem of noise resilience been carefully addressed, and an outstanding problem for theory is to identify the complexity threshold that must be crossed in order for CA to exhibit complex and robust emergent behavior. The constructions of Gacs [2, 30]—whose design calls to mind the genetic code of an elaborate artificial organism—provide extremely complicated starting points, but there is clearly much more to be done. We expect the NCA approach developed in the present work to be useful in this regard.
Acknowledgments:
We thank Shankar Balasubramanian, Sarang Gopalakrishnan, Vedika Khemani, Yaodong Li, Tibor Rakovsky, and Mike Zaletel for discussions and feedback. E.L. thanks Nathaniel Selub for discussions and collaboration on a related project. E.L. was supported by a Miller Research Fellowship.